OpenStudy (anonymous):

CAN SOMEONE HELP ME PLEASE IN GEOMETRY <3?

OpenStudy (anonymous):

A line that contains the points (2,-6) and (6,10). Another line graphed in the same coordinate plane contains the points (-4,9) and (8,6) Based on te slopes of these lines are they parallel, Perpendicular, or neither? Show all work.

OpenStudy (anonymous):

you have to find the slope of each line then compare them. when the slope is parallel they are the same. when the slope is perpindicular it's the negative reciprocal of the other slope. for example if one slope is 2 the perpindicular slope is -1/2

OpenStudy (anonymous):

That doesnt Hlep <3 im stupid xD

OpenStudy (anonymous):

To find the slope of each line we use y2-y1/x2-x1. We do this with our first line and get- 10-(-6)/6-2=16/4=4. This means our slope for the first line is 4. Now we do the same with the second line and get- 6-9/8-(-4)=-3/12=-1/4. That means our slope for the first line ie 4, and for the second -1/4. Now if they were parallel lines, the slopes would be exactly the same, which they aren't. If they are perpendicular, they would be the opposite negative. if you do the opposite of 4, you get 1/4 and the negative of that is -1/4 so that means your lines are perpendicular. -Hope that helps :)

OpenStudy (anonymous):

Omg thanks man your the best <3 Haha i think i may need your help more If you dont mind :)

OpenStudy (anonymous):

I'd be glad to help you out anytime if I can :) Sadly I g2g for now so yeah can't help for anymore. Just remember the slope formula is y2-y1/x2-x1. You have the two coords. of your line (x1,y1) & (x2,y2) You would just put them on that formula and that's how you would find your slope. :)

OpenStudy (anonymous):

ok thanks buddy <3 hold on im writing the info down xD

OpenStudy (anonymous):

K :)

OpenStudy (anonymous):

doing something for mom brb @Loujoelou @jordyn77

OpenStudy (anonymous):

Write the equation of the line that passes through the points (12.6) and (12,-4). Show all your work.

OpenStudy (anonymous):

@jordyn77 I need your help missy <3

OpenStudy (anonymous):

Write the equation of the line that passes through the points (12.6) and (12,-4). Show all your work.

OpenStudy (anonymous):

is that supposed to be (12,6)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You there ? lol

OpenStudy (anonymous):

lol yeah im figuring it lol lol

OpenStudy (anonymous):

oh sorry teacher

OpenStudy (anonymous):

you dont have to if you dont want too .. its ok :)

OpenStudy (anonymous):

lol im finishing calm down haha

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (anonymous):

did you really just write that on a paper and take a picture of it?

OpenStudy (anonymous):

lmao i just scanned it

OpenStudy (anonymous):

Awee your soo sweet thanks :)

OpenStudy (anonymous):

i did it for another problem too lol

OpenStudy (anonymous):

:,) Hold on i might need help with one more xD let me copy this down first :)

OpenStudy (anonymous):

lol ok, do you understand it?

OpenStudy (anonymous):

Yes :) Thanks to you

OpenStudy (anonymous):

lol okay.

OpenStudy (anonymous):

Write the equation of a line perpendicular to the line y=3x+9 and containing the points (6,-4). Show all work for credit <3

OpenStudy (anonymous):

;X thanks again for your help :)

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Lol thanks anyways :)

OpenStudy (anonymous):

lol no problem (:

OpenStudy (anonymous):

Hello @jordyn77 are you there?????

OpenStudy (anonymous):

lol yeah i sent you like three messages lol

OpenStudy (anonymous):

meee tooo? Something is wrong with mine :( i cant see anybodys fans or smart scores... </3

OpenStudy (anonymous):

imma restart my computer brb

OpenStudy (anonymous):

did you get my messages?

OpenStudy (anonymous):

^ @jordyn77

OpenStudy (anonymous):

yeah i answered.

OpenStudy (anonymous):

OMG I DIDNT GET IT -_____- Am i band or something???/ Like Whats wrong with my Thing

OpenStudy (anonymous):

im really getting aggravated ....

OpenStudy (anonymous):

im about to make a new one -_-

OpenStudy (anonymous):

lol go to chat for a second. thanks.

OpenStudy (anonymous):

im in the chat i dont see any messages ......

OpenStudy (anonymous):

okay well calm down for a few seconds kay thanks.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

lol change your picture just in case haha

OpenStudy (anonymous):

can you see anyone's fans and stuff?

OpenStudy (anonymous):

No i cant

OpenStudy (anonymous):

i cant see any fans or any Smart scores.. or any messages or anything just this convo and the stuff on the left

OpenStudy (anonymous):

okay, start a question on openstudy feedback

OpenStudy (anonymous):

hold on.

OpenStudy (anonymous):

okay start a question on OpenStudy Feedback not on here lol

OpenStudy (anonymous):

ok are you alive?